OpenStudy (anonymous):

Felipe drove 335 miles using 15 gallons of gas. At this rate, how many miles would he drive using 9 gallons of gas?

OpenStudy (anonymous):

m/ 9 = 335/15 -> m = ...?

OpenStudy (anonymous):

22.3

OpenStudy (anonymous):

Try again :/